Shift report 20010922 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Sep 22 2001 - 23:39:24 EDT

  • Next message: Apache: "Shift report 20010923 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Sep 22 23:39:23 2001, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: none
    Log book#9 pages 26 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    Machine status: Quelle machine!! RHIC was down for the whole shift
    QLI expert came in and was later joined by RF expert.
    Some beam circulated in the blue beam.
    
    Achievements: I used the new TPC pedestal "Inspector" XTPI installed 
    in pii4, great!
    I wish I could print those plots. TPM2 had some groups of pads
    with some weird spikes. See log book pg 27.
    I also looked at PMT and Si pedestals and there were some flags raised 
    for tiles3-1 and 4-5 as well as Si5-3.
    Then I decided to take drift velocity data every hour to see 
    how it changes with time. Some trend is clear in the 7 points I 
    collected.
    
    Detector problems: nothing new besides what I said above. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Sep 22 2001 - 23:39:47 EDT